OpenStudy (loser66):

\(f(x) =\begin{cases}x^rsin(1/x) ~~if~~x>0\\0~~if~~x\leq 0\end{cases}\) find values of r such that f(x) is everywhere differentiable. Please, help

OpenStudy (loser66):

@wio @freckles @mathmate @M4thM1nd

OpenStudy (loser66):

Suppose I know how to argue some steps and I get \(rx^{r-1}sin(1/x)+x^rlnx cos(1/x)=0\)

OpenStudy (loser66):

that is either x =0 or \(-\dfrac{r}{xlnx}= cot (1/x)\) then???

OpenStudy (kainui):

I think you've written \[\Large \frac{d}{dx} \frac{1}{x} \ne \ln x \\ \Large \frac{d}{dx} x^{-1} = -x^{-2}=\frac{-1}{x^2}\]

OpenStudy (loser66):

I don't get what you mean, how it relates to r?

OpenStudy (kainui):

When you did the chain rule after the product rule on taking the derivative here: \(rx^{r-1} \sin(1/x)+x^r \ln x \cos(1/x)=0\) I'm saying it should be \(rx^{r-1} \sin(1/x)+x^r (-x^{-2}) \cos(1/x)=0\)

OpenStudy (loser66):

oh yeah, I see my mistake

OpenStudy (loser66):

thanks a lot, let me redo, I may get something else.

OpenStudy (loser66):

@Kainui I got it, hehehe... thanks a lot.

OpenStudy (kainui):

Cool =D